OpenStudy (anonymous):

rationalize denominator : please help 1.) 4/ sqrt. 7p 2.) srqt.98/x 3.) 9/3 sqrt.y 4.) sqrt.7 / sqrt.5 +7

OpenStudy (anonymous):

1.) 4/7p−−√ 2.) 98−−√/x 3.) 9/3y√ 4.) 7√/5√+7

OpenStudy (anonymous):

\[1.) 4/\sqrt{7p}\]

OpenStudy (anonymous):

2.) \[\sqrt{98}/x\]

OpenStudy (anonymous):

The denominator is not supposed to have a square root. So when you rationalize you get rid of the square root in the denominator by multiplying the numerator and denominator by the square root, which is essentially multiplying by 1 so it doesnt change the value of the expression.

OpenStudy (anonymous):

3.) \[9/^{3}\sqrt{y}\]

OpenStudy (anonymous):

4.) \[\sqrt{7} / \sqrt{5} + 7\]

OpenStudy (dumbcow):

haha thats why its a cubed root, not 9/3

OpenStudy (anonymous):

yes i sent you the correction in the other post

OpenStudy (anonymous):

xavier : so i take out the sqrt ?

OpenStudy (anonymous):

LIke for 4/sqrt(7p). Multiply the numerator and denominator by sqrt(7p). So the denominator gets "rationalized" because sqrt(7p)*sqrt(7p) is simply 7p.

OpenStudy (anonymous):

so the answer would be : 4 sqr.7p ? or 4sqrt.7p/7p ?

OpenStudy (anonymous):

or ?

OpenStudy (anonymous):

4sqrt(7p)/7p

OpenStudy (anonymous):

and fot the second one ? would it be 7sqrt.2x ? or 7sqrt. 2/x ?

OpenStudy (anonymous):

\[7 \sqrt{2x} , 7 \sqrt{2/x}\]